OpenStudy (anonymous):

Simplify cos^2 Θ(1 + tan^2 Θ)

OpenStudy (math&ing001):

cos²Θ(1+tan²Θ) = cos²Θ + cos²Θ * tan²Θ = cos²Θ + sin²Θ ( tanΘ=sinΘ/cosΘ ) = 1

OpenStudy (anonymous):

\[\cos ^{\theta}(1+\frac{ \sin ^{2}(\theta) }{\cos ^{2}(\theta) })\]

OpenStudy (anonymous):

\[\cos ^{2}(\theta)+\sin ^{2}(\theta)=1\] This is a trig identity

OpenStudy (anonymous):

here are the answers: (1) cos2 Θ + sin2 Θ (2) cos2 Θ (3) 1 (4) −1

OpenStudy (anonymous):

according to the the two solutions posted by @math&ing001 and I what looks to be the correct answer?

OpenStudy (arbershabani97):

(3) 1 is the correct answer @math&ing001 is right

OpenStudy (anonymous):

i understand the distribution os²Θ(1+tan²Θ) = cos²Θ + cos²Θ * tan²Θ and I know that ( tanΘ=sinΘ/cosΘ ) How does this give me answer of 1?

OpenStudy (anonymous):

you have to know the trig identity \[\sin ^{2}(\theta)+\cos ^{2}(\theta)=1\] This is something you have to memorize

OpenStudy (arbershabani97):

http://www.heykiki.com/blog/wp-content/uploads/2012/12/a.gif here are some trig identities, that may help you understand them better

OpenStudy (anonymous):

i know this by heart sin2(θ)+cos2(θ)=1 here is what i need to understand, how does "cos²Θ + cos²Θ * tan²Θ" equals "cos²Θ + sin²Θ" I am not getting how the "tan²Θ" disappears cos²Θ(1+tan²Θ) = cos²Θ + cos²Θ * tan²Θ = cos²Θ + sin²Θ = 1

OpenStudy (anonymous):

http://www.amsi.org.au/teacher_modules/D5/D5g1.png here is an image demonstrating the trig identity

OpenStudy (anonymous):

\[\tan ^{2}(\theta)=\frac{ \sin ^{2}(\theta) }{ \cos ^{2}(\theta) }\]

OpenStudy (anonymous):

multiply tan^2(x) and cos^2(x) and you are left with just sin^2(x)

OpenStudy (anonymous):

got it :) THANK YOU ALL!